The serial numbers on $5 bills include three letters followed by seven digits. Assuming the digits are assigned at random, what is the probability that a serial number will contain,
(a)exactly two 5s
(b) at least four 5's
(c) all 5s?

Answer by Boreal(15235)

10 digits so for two 5s, it would be 7C2, the number of different places the 5s can occur, multiplied by 0.1^2*0.9^5=0.1240
at least four: do 0, 1,2,3 and subtract from 1. For 0, it would be.9^7=0.4783
For 1 it would be 7*.1*.9^6=0.3720
For 2 it would be 0.1240 from above
For 3 it would be 35 (7C3)*.1^3*.9^4=0.0230
That sum is 0.9973 so the probability of at least four 5s is 0.0027
Can check by doing 4, which is 0.00255 and 5 which would be 0.00017.
Probability all 5s is .1^7, or 1 x 10^(-7)